### v1

# Harnessing Consistency for Robust Test-Time LLM Ensemble

Abstract
[…]
Harnessing Consistency for Robust Test-Time LLM Ensemble
Zhichen Zeng1  Qi Yu1  Xiao Lin1  Ruizhong Qiu1  Xuying Ning1 Tianxin Wei1Yuchen Yan1Jingrui He1Hanghang Tong1 1University of Illinois Urbana-Champaign {zhichenz, htong}@illinois.edu
1 Introduction
Large language models (LLMs) Brown et al. (2020); Team et al. (2023); Touvron et al. (2023); Achiam et al. (2023); Guo et al. (2025) have demonstrated remarkable performance in natural language processing tasks. Due to the difference in model architectures, training algorithms, and datasets, different LLMs expertize in different areas, and it is important to ensemble various LLMs to integrate their complementary knowledge Yao et al. (2024); Abdulaal et al. ; Huang et al. (2024).
Extensive efforts on test-time LLM ensemble can be broadly categorized into two categories: token-level and response-level ensemble. Token-level ensemble aligns and fuses the token probabilities of different LLMs at each decoding step Yu et al. (2024); Yao et al. (2024); Huang et al. (2024); Xu et al. (2024c), enabling fine-grained real-time correction for each token generation. Response-level ensemble offers a more coarse-level ensemble by selecting either a complete response Jiang et al. (2023b); Lv et al. (2024a); Tekin et al. (2024) or a span Xu et al. (2024b); Lv et al. (2024b); Liu et al. (2024) from candidate outputs. Despite their success in improving ensemble quality, existing approaches largely overlook ensemble robustness against noisy or erroneous signals. For instance, incorrect token alignments can result in faulty probability fusion, while errors in model predictions may further compromise the correctness of ensemble outputs. Therefore, it is crucial for ensemble methods to detect and mitigate such potential errors during inference to ensure reliable performance.
Figure 1: Motivation of CoRE. Left: vanilla ensemble yields an incorrect prediction because the token _r, misaligned from the assistant token rine, dominates the ensemble. Right: CoRE penalizes inconsistent tokens, rendering correct prediction r.
To bridge this gap, we propose a plug-and-play technique named CoRE to enhance both robustness and performance of LLM ensemble. Our preliminary observations reveal that ensemble failures are closely tied to large discrepancies between the token distributions of different LLMs. Motivated by this, we propose to harness both token and model consistency to achieve robust ensemble. At the token level, significant disparities in output probabilities for a specific token indicate fine-grained uncertainty, often stemming from token space misalignment. To address this, we introduce token consistency, which measures the disparity between each model’s token probability and a reference probability, serving as a low-pass filter to amplify reliable tokens and suppress inconsistent ones. At the model level, large divergence among full token distributions reflects model conflicts. To capture this, we define model consistency that promotes model outputs with high self-confidence and low divergence from peers, thereby strengthening consistent models and downweighting unreliable ones. A key advantage of the proposed CoRE is that it is orthogonal to various token-level ensemble methods, and thus can be seamlessly integrated with no additional inference cost.
We summarize the main contributions as follows
  • We are the first to investigate the robustness issue in LLM ensemble.
  • We assess consistency at both the token and model levels to enhance ensemble performance and robustness. This method can be seamlessly integrated with various LLM ensemble strategies to enable test-time ensemble correction with no additional cost.
  • We conduct extensive experiments across diverse benchmark tasks, model combinations, and ensemble methods. As shown in Figure 2, CoRE consistently improves baseline ensemble methods, achieving an average performance gain of 1.3% and 2.8% on Top-2 and Top-3 model ensemble, respectively.
Figure 2: Ensemble performance across six benchmarks. We report the average performance of different ensemble methods with (CoRE) and without (Vanilla) CoRE.
2 Related Works
2.1 Test-time LLM Ensemble
Ensembling multiple large language models (LLMs) at test time offers a practical way to harness their diverse strengths and mitigate individual weaknesses. Existing methods can be broadly categorized into token and response level ensembles.
Token‑level ensemble fuses next-token predictions across models at each decoding step, enabling fine-grained correction during generation. Early works Fu et al. (2023); Wan et al. (2024); Mavromatis et al. (2024) align token sequences via minimum edit distance, capturing structural difference but incurring high computational cost. GAC Yu et al. (2024) and UniTE Yao et al. (2024) bridge disparate vocabularies by aligning tokens through exact or prefix matches in text space. Recent works, such as DeepEn Huang et al. (2024) and EVA Xu et al. (2024c), learn projection functions that map heterogeneous token distributions into a shared representation space, enabling direct output fusion.
Response-level Ensemble selects the most promising response from the multiple outputs to ensemble LLMs at a coarser-grained level. One line of work selects or synthesizes a full response among model outputs by either training-free approaches, such as perplexity scoring or majority voting Jiang et al. (2023b); Lv et al. (2024a); Tekin et al. (2024), or training-based approaches that learn to rank responses Si et al. (2023). Another line of work Xu et al. (2024b); Lv et al. (2024b); Liu et al. (2024) proposes ensembling at the span level, aiming to strike a balance between fine-grained correction and context-aware decision-making.
Figure 3: Overview of CoRE. Token probabilities from different models are first aligned through token mapping, and a reference probability distribution is constructed for consistency evaluation. Token consistency mitigates the influence of inconsistent tokens, while model consistency highlights reliable and self-consistent models. Combining both yields a robust and improved ensemble prediction.
2.2 Model Consistency
Recent research has explored consistency-based strategies to boost LLM performance by either internal self-consistency or cross-model agreement.
Self-Consistency enhances answer reliability by aggregating diverse reasoning paths from a single model via frequency Wang et al. (2022); Li et al. (2024a); Aggarwal et al. (2023), entropy Kadavath et al. (2022); Lin et al. (2023); Kang et al. (2025), or confidence signals Chen et al. (2023); Taubenfeld et al. (2025), but it incurs notable computational overhead due to repeated sampling.
Multi-model Consistency combines outputs from different LLMs through majority voting Trad and Chehab (2024); Niimi (2025) or collaborative reasoning Wang et al. (2024); Liang et al. (2023); Li et al. (2024b). Despite improving robustness, most focus on response-level agreement, overlooking fine-grained token-level consistency essential for detecting subtle errors in real time.
3 Methodology
In this section, we present our proposed CoRE for trustworthy LLM ensemble. We begin by introducing the notations and formally defining the LLM ensemble problem in Section 3.1. Then, in Section 3.2, we reveal a strong correlation between ensemble performance and consistency scores at both the token and model levels. Motivated by these findings, we introduce our token-level consistency measure in Section 3.3 and our distribution-level consistency measure in Section 3.4, both designed to improve ensemble performance and reliability.
(a) Token Probability Disparity.
(b) Entropy of Token Probability.
(c) Sum of Token Consistency.
Figure 4: Observations. (a) Token probability disparity: Aligned tokens (blue) exhibit smaller disparities than misaligned ones (orange), indicating higher consistency. (b) Entropy of token probability: Correct answers (blue) exhibit lower entropy than wrong ones (orange), indicating higher confidence. (c) Sum of token consistency: Correct answers (blue) exhibit higher sum of token consistency score than wrong ones (orange).
3.1 Problem Formulation
Formally, we are given a main model with vocabulary set $\mathcal{V}_{\text{main}}$ and $N$ assistant models with vocabulary set $\mathcal{V}_{\text{assist}_{i}}$ with $i\in\{1,2,...,N\}$. We denote the predicted token distributions of main and assistant models as $\bm{p}_{\text{main}}$ and $\bm{p}_{\text{assist}_{i}}$, respectively. LLM ensemble first learns a token alignment matrix $\bm{A}_{i}\in\mathbb{R}^{|\mathcal{V}_{\text{assist}_{i}}|\times|\mathcal{V}_{\text{main}}|}$ between the main model and the $i$-th assistant model, where $\bm{A}_{i}(v,u)=1$ indicates that token $v\in\mathcal{V}_{\text{main}}$ is aligned with token $u\in\mathcal{V}_{\text{assist}_{i}}$. Token prediction probabilities can be further projected into the main model token space via $\bm{p}_{\text{assist}_{i}}\bm{A}_{i}\in\mathbb{R}^{|\mathcal{V}_{\text{main}}|}$. For simplicity, we denote the aligned probability as $\tilde{\bm{p}}_{\text{assist}_{i}}=\bm{p}_{\text{assist}_{i}}\bm{A}_{i}$. The aligned probabilities can be further ensembled via
$\bm{p}_{\text{ens}}=w_{\text{main}}\bm{p}_{\text{main}}+\sum_{i=1}^{N}w_{\text{assist}_{i}}\tilde{\bm{p}}_{\text{assist}_{i}}$
where $\left[w_{\text{main}},w_{\text{assist}_{1}},...,w_{\text{assist}_{N}}\right]\in\Delta^{N+1}$ are the normalized model weights. For the special case where all models are assigned uniform weights, we denote the resulting ensembled probability distribution as $\bm{p}^{*}$, that is
$\bm{p}^{*}=\frac{1}{N+1}\left(\bm{p}_{\text{main}}+\sum_{i=1}^{N}\tilde{\bm{p}}_{\text{assist}_{i}}\right),$ (1)
In this paper, we use the above average probability as the reference $\bm{p}^{*}$ for consistency computation.
3.2 Observations
We first study the distinct patterns in both token and model levels, which signal the inherent factors affecting the ensemble performance in Figure 4.
First, we examine how token disparity, measure by the difference between the aligned probability and the reference probability on a specific token $v$, i.e., $\bm{\delta}_{i}(v)=|\tilde{\bm{p}}_{\text{assist}_{i}}(v)-\bm{p}^{*}(v)|$, reflects the alignment correctness. As shown in Figure 4(a), aligned tokens concentrate at lower disparities, while misaligned tokens exhibit a heavier right tail. A one-sided test of $H_{0}:\mathbb{E}_{v\in\text{misalign}}\bm{\delta}_{i}(v)\leq\mathbb{E}_{v\in\text{align}}\bm{\delta}_{i}(v)$ yields a $p$-value of $2.7\times 10^{-44}$, indicating that aligned tokens indeed have smaller disparity. We summarize this in Observation 1.
Observation 1 (Token Consistency).
Large token probability disparity signals token misalignment.
For example, if token $\texttt{a}\in\mathcal{V}_{\text{assist}_{i}}$ is misaligned to token $\texttt{b}\in\mathcal{V}_{\text{main}}$, the difference in their probabilities $|\bm{p}_{\text{assist}_{i}}(\texttt{a})-\bm{p}^{*}(\texttt{b})|=|\tilde{\bm{p}}_{\text{assist}_{i}}(\texttt{b})-\bm{p}^{*}(\texttt{b})|$ is expected to be large.
Secondly, we evaluate how model confidence, measured by the entropy of the token probability $H(\tilde{\bm{p}}_{\text{assist}_{i}})$, relates to the answer correctness. As shown in Figure 4(b), correct answers exhibit lower entropy than wrong ones. A one-sided test of $H_{0}:\mathbb{E}[H(\tilde{\bm{p}})|\text{wrong}]\leq\mathbb{E}[H(\tilde{\bm{p}})|\text{correct}]$ yields a $p$-value of $7.7\times 10^{-108}$, confirming that wrong answers have larger entropy than correct ones. We summarize this into Observation 2.
Observation 2 (Model Confidence).
Small model confidence measure by the entropy of token probability signals correct answer.
Thirdly, we evaluate how model consistency, measured by the sum of RBF-transformed token disparities, i.e., $C_{\text{assist}_{i}}(v)=\allowbreak\sum_{v\in\mathcal{V}_{\text{main}}}\exp(-\bm{\delta}_{\text{assist}_{i}}(v)/\sigma)$, varies between correct and wrong answers. As shown in Figure 4(c), correct answers exhibit higher model consistency than wrong ones. A one-sided test of $H_{0}:\mathbb{E}[C_{\text{assist}_{i}}|\text{wrong}]\geq\mathbb{E}[C_{\text{assist}_{i}}|\text{correct}]$ yields a $p$-value of $5.0\times 10^{-222}$, indicating that correct answers have greater consistency. We summarize this into Observation 3.
Observation 3 (Model Consistency).
Large model consistency measured by the sum of RBF-transformed token disparity signals correct answer.
Motivated by these observations, we propose to harness both token and model consistency for robust LLM ensemble.
3.3 Token Consistency
We first introduce our proposed token consistency as a fine-grained measure. A key challenge in LLM ensemble is the disparate token spaces due to different tokenization schemes adopted by different LLMs. Despite extensive efforts in aligning token spaces, they inevitably encounter alignment errors that impair ensemble performance.
To address this limitation, it is essential to distinguish the misaligned tokens and rectify their contributions. Motivated by Observation 1, where misaligned tokens exhibit large probability disparities, we propose token consistency as a low-pass filter to downweight their influence. Adopting the average distribution in Eq. (1) as the reference probability, we quantify the $i$-th assist model’s token consistency $\bm{s}^{\text{t}}_{\text{assist}_{i}}$ as follows
$\bm{s}^{\text{t}}_{\text{assist}_{i}}=f(\bm{\delta}_{i})\in\mathbb{R}^{|\mathcal{V}_{\text{main}}|},$ (2)
$\text{where }\bm{\delta}_{i}=|\tilde{\bm{p}}_{\text{assist}_{i}}-\bm{p}^{*}|\in\mathbb{R}^{|\mathcal{V}_{\text{main}}|},$
where different functions, e.g., RBF kernel $f_{\text{rbf}}(\bm{\delta})=\exp(-\bm{\delta}/\sigma)$, power function $f_{\text{pow}}(\bm{\delta})=\alpha(1-\bm{\delta})^{\beta}$, and sigmoid function $f_{\text{sig}}(\bm{\delta})=1-\text{Sigmoid}(\gamma(\bm{\delta}_{i}-0.5))$, can be adopted. Intuitively, large $\bm{\delta}$ induces small token consistency that signals large inconsistency with the reference probability. By multiplying token consistency with the aligned distribution, i.e., $\bm{s}^{\text{t}}_{\text{assist}_{i}}\odot\tilde{\bm{p}}_{\text{assist}_{i}}$, token consistency acts as a low-pass filter: penalizing inconsistent tokens with large disparities while promoting consistent tokens that are widely agreed upon.
3.4 Model Consistency
In addition to fine-grained token consistency, it is crucial to quantify the trustworthiness of the model. Prior works typically assign model weights based on heuristics, such as uniform weighting Yao et al. (2024); Yu et al. (2024); Xu et al. (2024c) or prior knowledge Huang et al. (2024), or self-confidence metrics like perplexity Mavromatis et al. (2024); Liu et al. (2024). However, the inter-model consistency remains largely underexplored.
Motivated by Observation 2 and 3, we argue that models exhibiting both high inter-model consistency and strong self-confidence should be prioritized. To capture this, we define the model consistency by aggregating token consistency over the main token space, and regularizing it by the entropy of the output distribution serving as a proxy for confidence. Formally, the model consistency of an assistant model is given by:
$s^{\text{m}}_{\text{assist}_{i}}=\frac{\sum_{v\in\mathcal{V}_{\text{main}}}\bm{s}^{\text{t}}_{\text{assist}_{i}}(v)}{H(\tilde{\bm{p}}_{\text{assist}_{i}})}\in\mathbb{R},$ (3)
where $H(\cdot)$ denotes the entropy of a distribution. Here, the numerator rewards agreement with the reference model, while the denominator penalizes high uncertainty, thus favoring outputs that are both consistent and confident. A similar definition applies for the main model, denoted as $s^{\text{m}}_{\text{main}}$.
By using token consistency as a token-wise filter and model consistency as model-level weights, the final ensembled distribution is computed as:
$\bm{p}_{\text{ens}}=s^{\text{m}}_{\text{main}}\bm{p}_{\text{main}}+\sum_{i=1}^{N}s^{\text{m}}_{\text{assist}_{i}}\bm{s}^{\text{t}}_{\text{assist}_{i}}\odot\tilde{\bm{p}}_{\text{assist}_{i}},$ (4)
where $s^{\text{m}}_{\text{main}},s^{\text{m}}_{\text{assist}_{1}},\dots,s^{\text{m}}_{\text{assist}_{N}}]\in\Delta^{N+1}$ are the normalized model consistency serving as the model weights. Note that we apply token consistency only to assistant models to mitigate potential misalignment, as token misalignment arises solely when mapping tokens from assistant models to the main model’s token space.
4 Experiments
We carry out extensive experiments to answer the following research questions:
  • RQ1: How does CoRE enhance ensemble performance? (Section 4.2)
  • RQ2: To what extent does CoRE enhance ensemble robustness? (Section 4.3)
  • RQ3: What are the respective roles of token and model consistency, and how do their designs affect performance? (Section 4.4)
  • RQ4: How does the performance scale with more models w/ and w/o CoRE? (Section 4.4)
4.1 Experiment Setup
Base Models.
We conduct our experiments on the following widely used models, including Llama-3-8B-Instruct Dubey et al. (2024), Mistral-7B-Instruct-v0.1 Jiang et al. (2023a), Qwen2.5-3b-Instruct Team (2024), InternLM2.5-7b-Chat Team (2023) and openchat-3.5-0106 Wang et al. (2023).
Ensemble Methods.
We consider four baseline methods that align token spaces for LLM ensemble for benchmark evaluation, including:
  • MinED Fu et al. (2023); Mavromatis et al. (2024) searches for textually closest tokens based on minimum edit distance.
  • GaC Yu et al. (2024) merges disparate token spaces into a union token space for ensemble;
  • UniTE Yao et al. (2024) utilizes tokenizer to map tokens to their prefix counterparts;
  • EVA Xu et al. (2024c) learns a mapping by aligning overlapping token embeddings.
Note that our proposed CoRE mainly operates on the token space and is not specifically designed for methods like DeePEn Huang et al. (2024) that ensemble in the latent embedding space. While being outside our main scope, we nonetheless conduct a separate analysis by slightly adapting CoRE to integrate with DeePEn, in order to explore its potential benefits in this alternative setting.
Method GSM8K PIQA SAMSum TriviaQA NQ MMLU Average
Top-2 Top-3 Top-2 Top-3 Top-2 Top-3 Top-2 Top-3 Top-2 Top-3 Top-2 Top-3 Top-2 Top-3
MinED
Vanilla 79.51 81.65 85.47 82.32 43.79 42.18 67.50 59.63 25.07 22.22 68.38 66.66 61.62 59.11
CoRE 82.41 83.78 85.75 87.49 49.25 49.55 72.17 72.85 29.83 29.61 68.42 69.77 64.64 65.51
[…]
UniTE
Vanilla 81.26 82.79 85.64 86.89 48.03 47.99 70.93 64.90 28.17 25.84 68.61 69.56 63.78 63.00
CoRE 82.71 83.70 85.96 87.38 49.41 49.77 73.41 73.61 30.39 29.94 68.63 69.91 65.09 65.72
[…]
EVA
Vanilla 82.09 83.02 87.81 87.43 48.16 47.83 73.47 73.02 30.00 29.09 70.62 71.07 65.36 65.24
CoRE 83.40 83.24 87.81 87.76 49.43 50.27 73.13 73.23 29.53 29.09 70.62 70.91 65.65 65.75
[…]
GaC
Vanilla 80.74 83.02 85.67 86.89 48.08 47.63 73.67 71.07 29.61 27.15 68.61 69.56 64.40 64.22
CoRE 82.49 84.00 85.75 87.32 49.41 49.73 73.72 73.97 30.22 29.81 68.63 69.96 65.04 65.80
[…]
Table 1: Benchmark results. We report ensemble performance using the Top-2 and Top-3 base models on each dataset, with (CoRE) and without (Vanilla) our method. The $\Delta$ rows report the performance gain from applying CoRE, with blue cells indicating improvement and red cells indicating degradation.
Model GSM8K PIQA SAMSum TriviaQA NQ MMLU
[…]
Mistral7b 41.55 $62 \cellcolor$[HTML]FFF2CC44.80 52.47 15.01 52.17
[…]
Table 2: Base model performance. We use Blue, Yellow and Red to denote Top-1, Top-2 and Top-3 models.
Datasets and Metrics.
We evaluate six benchmarks covering four different categories, including: (1) Reasoning: GSM8K Cobbe et al. (2021) (4-shot with CoT) covering grade school math problems, and PIQA Bisk et al. (2020) (0-shot) with commonsense reasoning choice problems; (2) Summarization: SAMSum Gliwa et al. (2019) (0-shot) on dialogue summarization; (3) Knowledge: TriviaQA Joshi et al. (2017) (5-shot) and NaturalQuestions (NQ) Kwiatkowski et al. (2019) (5-shot); (4) Comprehensive Examination: MMLU Hendrycks et al. (2009) (5-shot) covering 57 subjects that humans typically learn. We adopt Exact Match for PIQA, TriviaQA, NQ and MMLU, Accuracy for GSM8K, and Rouge-1 score for SAMSum.
Experiment Pipeline.
We first evaluate the performance of base models on each dataset, and then select the Top-2 and Top-3 models for benchmark ensemble evaluation. We adopt RBF function as the default consistency score for benchmark results, and set $\sigma=0.5$. All experiments are conducted on NVIDIA A100 80GB GPUs.
4.2 Benchmark Results
We present the ensemble results in Table 1 and base models’ performance in Table 2, from which we draw the following observations:
(1) CoRE achieves consistent improvements on different methods, datasets, and base model combinations. Specifically, on reasoning datasets (GSM8K, PIQA), CoRE enhances vanilla methods by 1.01 on Top-2 and 1.33 on Top-3 ensemble. On the summarization dataset (SAMSum), CoRE improves vanilla baselines by an average of 2.35 and 3.42 on Top-2 and Top-3 ensembles, respectively. For knowledge-intensive datasets (TriviaQA, NQ), it yields average gains of 1.75 (Top-2) and 4.90 (Top-3). On the comprehensive exam benchmark (MMLU), CoRE provides smaller but consistent improvements of 0.03 (Top-2) and 0.94 (Top-3).
(2) CoRE achieves more stable ensemble. As more LLMs are included, baseline ensembles may suffer from negative ensemble, where performance degrades compared to the best single model. In contrast, augmenting with CoRE leads to robust and consistent improvements. Notably, CoRE successfully mitigates 17 negative ensemble cases encountered by the baseline ensemble methods.
Though not directly compatible with DeePEn, which performs ensemble in the latent embedding space rather than the token space, we adapt CoRE to compute consistency over token embeddings and report the results in Table 3. Though less pronounced than token-space baselines, augmenting with CoRE still yields consistent improvements. We attribute this to two main factors: (1) DeePEn constructs its latent space using overlapping tokens that are identical, which inherently reduces token misalignment, and (2) the learned latent space already promotes cross-model agreement by design. Interestingly, though some vanilla baselines, e.g., MinED and UniTE, may underperform DeePEn, they achieve better performance when augmented with CoRE, validating the effectiveness of CoRE.
Method PIQA NQ MMLU Average
Top-2 Top-3 Top-2 Top-3 Top-2 Top-3 Top-2 Top-3
Vanilla 87.52 87.35 28.03 29.23 70.56 70.49 62.04 62.36
CoRE 87.74 87.35 28.14 29.53 70.57 70.5 62.15 62.46
[…]
Table 3: Ensemble performance with DeePEn.
(a) Token noises.
(b) Probability noises.
Figure 5: Robustness against noises. CoRE maintains stable performance with minimal degradation under noises.
Figure 6: Robustness against large performance gap. We ensemble the best and worst performing models on NQ and TriviaQA. Numbers above the bars indicate the performance gains achieved by CoRE.
4.3 Robustness Results
4.3.1 Robustness against Noises
We first evaluate ensemble performance against noises. We consider two types of noises, including: (1) alignment noise, where 5%, 10%, 15% and 20% of the rows in the token mapping matrix are perturbed, and (2) probability noise, where Gaussian noise with standard deviations of 0.05, 0.10, 0.15, and 0.20 is added to the token probabilities. The results are shown in Figures 5.
Under both noise types, vanilla ensemble methods are highly sensitive, with average performance drops of 4.25 and 2.60 points as noise ratios increase from 0 to 0.2 for token and probability noise, respectively. In contrast, CoRE demonstrates strong robustness, exhibiting only marginal degradation of 0.38 and 0.49 under the same conditions. Notably, CoRE is particularly effective in mitigating the impact of token alignment noise. This improvement can be attributed to the token consistency mechanism that identifies and rectifies misaligned tokens, thereby preserving ensemble accuracy even under severe perturbations.
4.3.2 Robustness against Performance Gap
Previous methods Yao et al. (2024) may encounter negative ensemble, i.e., model performance drops after ensemble, when facing significant performance discrepancies among base models. To evaluate CoRE’s robustness against performance gaps, we ensemble the best and worst performing models on NQ and TriviaQA with and without CoRE, and the results are shown in Figure 6.
It is shown that CoRE consistently improves all baseline ensemble methods on two datasets. In particular, for vanilla methods that suffer severe degradation, e.g., MinED, UniTE, and GaC, augmenting with CoRE yields substantial average gains of +5.66 on NQ and +9.42 on TriviaQA. Moreover, while vanilla methods exhibit widely varying performance, their performance with CoRE converges to comparable levels, highlighting CoRE’s ability to mitigate performance gaps and deliver stable improvements across diverse ensemble strategies.
4.4 Studies
4.4.1 Ablation Study
We conduct ablation studies on token and model consistency on NQ and TriviaQA, and the results are shwon in Figure 7. A clear trend emerges: CoRE achieves the best performance, followed by applying only token consistency $\bm{s}^{t}$, then only model consistency $s^{m}$, while the vanilla ensemble performs the worst. This indicates that both consistency components contribute positively to the performance: token consistency enhances fine-grained alignment across tokens, model consistency improves global agreement among models, and their joint integration yields the best overall gains.
Figure 7: Ablation study. Both token and model consistency benefit the ensemble performance.
4.4.2 Scaling Results
We evaluate how CoRE enhances ensemble scalability as more LLMs are incorporated. As shown in Figure 8, vanilla ensembles suffer from negative ensembling, with performance often degrading as more models are added, whereas CoRE enables stable scaling and consistently outperforms the best single model. This performance degradation of the vanilla ensemble arises because increased model diversity introduces conflicting signals that, without consistency control, overwhelm the ensemble. Furthermore, token space alignment becomes more challenging when more LLMs are involved, and being unaware of the potential misalignment can lead to unstable and even counterproductive ensembling. However, CoRE enforces consistency and filters out unreliable outputs, transforming diversity into complementary information rather than noise, and thereby enabling robust and effective scaling.
Figure 8: Scaling results on TriviaQA. Left: Vanilla ensemble methods suffer from negative ensembling, with performance degrading as more models are added. Right: CoRE enables stable scaling, consistently outperforming the best single model across ensemble sizes.
4.4.3 Consistency Score Functions
We explore alternative designs for consistency score to evaluate the generalization of CoRE. Specifically, we consider three consistency functions: CoRE-RBF with RBF kernel $f_{\text{rbf}}(\bm{\delta})=\exp(-\bm{\delta}/\sigma)$, CoRE-pow with power function $f_{\text{pow}}(\bm{\delta})=\alpha(1-\bm{\delta})^{\beta}$, and CoRE-sig with sigmoid function $f_{\text{sig}}(\bm{\delta})=1-\text{Sigmoid}(\gamma(\bm{\delta}-0.5))$. Experiments are conducted on the SAMSum dataset, with results presented in Figure 9.
In general, different consistency score designs consistently enhance ensemble methods, demonstrating the broad applicability of CoRE. On average, CoRE-Pow achieves the best performance with a 2.76 gain, followed by CoRE-RBF (2.35) and CoRE-Sigmoid (2.06). The performance differences stem from how each function maps token disparity $\bm{\delta}$ to consistency weights. RBF sharply favors well-aligned tokens, offering stable but conservative gains. Power applies smoother decay, better balancing selectivity and inclusiveness, thus achieving the highest gain. Sigmoid filters noise via soft thresholding but its limited smoothness dampens overall improvement.
4.4.4 Case Study
Example 1 Question: what does the adrenal gland produce that is necessary for the sympathetic nervous system to function? Gold Answer: epinephrine OpenChat Response: adrenaline
InternLM Response: epinephrine and norepinephrine
Vanilla Response: epineph rine
CoRE Response: epinephrine
Figure 9: Ensemble performance with different consistency score functions on SAMSum. Numbers above the bars indicate the performance gains achieved.
Figure 10: The Top-5 next token probability and the corresponding token consistency $\bm{s}^{\text{t}}_{\text{assist}}$ given generated text ’epineph’ in Example 1.
We present a case study on ensembling OpenChat3.5 and InternLM2.5 offering an intuitive illustration of how CoRE operates. As shown in Example 1 and Figure 10, both individual models and vanilla ensemble fail to generate the correct answer, whereas augmenting with CoRE yields the correct response. The key difference lies in the fourth token following the generated text “epineph”: the vanilla ensemble predicts _r, while CoRE correctly predicts r. Analyzing the token probabilities of InternLM2.5 before and after token mapping reveals that the correct token rine is incorrectly aligned to _r, leading the vanilla ensemble to an erroneous fusion. In contrast, CoRE identifies the misaligned token _r given its low consistency and downweights its influence, thereby penalizing unreliable tokens and models to produce the correct output. More case studies are provided in Appendix B.
5 Conclusion
In this paper, we study the robustness of LLM ensembles, a critical yet underexplored aspect in prior work. Our analysis shows that ensemble failures often arise from inconsistencies at both the token and model levels. To address this, we introduce CoRE, a lightweight and plug-and-play framework that enforces consistency across multiple dimensions without incurring additional inference cost. Extensive experiments show that CoRE significantly enhances both the performance and robustness of existing ensemble methods. Our findings highlight consistency as a key principle for building reliable LLM ensembles and open new directions for robustness-oriented ensemble in future research.
Appendix
Appendix A Experiments
A.1 Examples of Token Misalignment
We provide intuitive examples of token misalignment caused by different ensemble methods in Table 4, validating that existing ensemble method may generate suboptimal or even irrelevant token mappings. Capable of handling token misalignment, CoRE consistently improves the performance of existing ensemble methods.
Table 4: Examples of misaligned tokens from source LLM (OpenChat) to mapped LLM (InternLM).
Method Source Token Mapped Token
MinED ’riv’ ’_IV’
UniTE ’_ind’ ’indows’
EVA ’equation’ ’align’
A.2 Adapting CoRE to DeePEn
To adapt CoRE for DeePEn that ensembles models in a latent embedding space, we normalize the unified token embeddings and transform them into probability distributions which CoRE can operate on before ensemble. This adaptation is natural because DeePEn, rather than projecting tokens of assist LLMs into the vocabulary space of the main LLM, essentially projects tokens of ensembled LLMs into a joint latent embeddings space on top of common tokens across different LLMs. By interpreting the normalized embeddings as token probability vectors, CoRE can be integrated seamlessly into embedding-based ensemble methods such as DeePEn.
Appendix B Additional Studies
B.1 Observation on Model Consistency
In addition to observations in 3.2, we directly evaluate how model score, measured by the quotient of model consistency and confidence, i.e., $s^{\text{m}}_{\text{assist}_{i}}=\sum_{v\in V_{\text{main}}}{\bm{s}^{\text{t}}_{\text{assist}_{i}}(v)/H(\tilde{\bm{p}}_{\text{assist}_{i}})}$, indicates the correctness of responses. As shown in Figure 11, correct answers exhibit higher model score than the wrong ones. A one-sided test of $H_{0}:\mathbb{E}[s^{\text{m}}_{\text{assist}_{i}}|\text{wrong}]\geq\mathbb{E}[s^{\text{m}}_{\text{assist}_{i}}|\text{correct}]$ yields a $p$-value of $2.5\times 10^{-79}$, confirming that correct answers have larger model scores than the wrong ones. We summarize this into Observation 4, which further validates our design of the model score.
Observation 4 (Model Score).
Large model score measured by the quotient of model consistency and confidence signals correct answer.
Figure 11: Model score visualization.
B.2 Additional Case Studies
We present additional case studies on the Comprehensive Examination dataset (MMLU), Summarization dataset (SAMSum), and Reasoning dataset (GSM8K) to provide a more intuitive understanding of how CoRE operates. As illustrated in Figure 12, correct answers are highlighted in green and incorrect ones in red. CoRE successfully produces the correct answers where both individual base models and the vanilla ensemble fail.
Example 2 (Comprehensive Examination) Question: A person weighs 62 kg. Their drug dose is 15 mg/kg. How many grams is their dose? Choose one answer from the following: A. 930  B. 93  C. 9.3  D. 0.93 Gold Answer: D InternLM Response: C
Qwen Response: C
Vanilla Response: C
CoRE Response: D
Example 3 (Summarization) Question: Generate a concise summary for the chat: Kelly: Should I put on the red dress? Meghan: oh yes!!!! Mary: I’$l$ wear a red dress too, and bright, red lipstick Kelly: bloody Mary! Mary: hahahaha Gold Answer: Kelly and Mary will wear red dresses. Mary will wear a red lipstick too. OpenChat Response: Kelly, Meghan, and Mary all decide to wear red dresses and bright red lipstick. (Rouge Score=0.55) Mistral Response: Kelly is considering wearing a red dress and Mary suggests she wear red lipstick to go with it. (Rouge Score=0.64) Vanilla Response: Kelly is considering wearing a red dress and Mary suggests wearing a red dress and red lipstick. Kelly responds with "bloody Mary". (Rouge Score=0.55) CoRE Response: Kelly is considering wearing a red dress and Mary will also wear a red dress and red lipstick. (Rouge Score=0.73)
Example 4 (Reasoning) Question: John plans to sell all his toys and use the money to buy video games. He has 13 lego sets and he sells them for \$15 each. He ends up buying 8 video games for \$20 each and has \$5 left. How many lego sets does he still have? Gold Answer: He spent \$160 on video games because 8 x 20 = 160 He earned \$165 from his sale because 160 + 5 = 165 He sold 11 lego sets because 11 x 15 = 165 He has 2 sets left because 13 - 11 = 2 Therefore, the answer is 2 InternLM Response: John sells 13 lego sets for \$15 each, so he earns 13*15 = \$195 He buys 8 video games for \$20 each, so he spends 8*20 = \$160 He has \$5 left, so the total amount of money he earned is \$160 + \$5 = \$165 Since he earned \$195 from selling his lego sets, the number of lego sets he still has is \$195 - \$165 = \$30 The answer is 30 ✗ OpenChat Response: John sells his 13 lego sets for \$15 each, so he makes 13*\$15=\$195 He uses this money to buy 8 video games for \$20 each, so he spends 8*\$20=\$160 He has \$5 left, so he spent \$195-\$160-\$5=\$30 on lego sets Since he sells each lego set for \$15, he sells 30/15=2 lego sets He started with 13 lego sets and sold 2, so he still has 13-2=11 lego sets The answer is 11 ✗ Vanilla Response: John sells 13 lego sets for \$15 each, so he earns 13*15 = \$195 He buys 8 video games for \$20 each, so he spends 8*20 = \$160 He has \$5 left, so he spent a total of 195-5 = \$190 on lego sets Since he sold 13 lego sets for \$15 each, he must have had 190/15 = 13 lego sets The answer is 13 ✗ CoRE Response: John sells 13 lego sets for \$15 each, so he earns 13*15 = \$195 from selling the lego sets. He buys 8 video games for \$20 each, so he spends 8*20 = \$160 on video games. He has \$5 left, so he spent a total of 195-160-5 = \$30 on other things. Since each lego set is sold for \$15, he still has 30/15 = 2 lego sets left. The answer is 2 ✓
Figure 12: Additional case study across comprehensive examination, summarization, and reasoning benchmarks. Correct and incorrect responses are highlighted in green and red, respectively.
Appendix C Datasets
GSM8K Cobbe et al. (2021) is a multi-step arithmetic reasoning dataset consisting of 1,319 linguistically diverse grade-school math word problems authored by humans. Each problem requires explicit intermediate reasoning to reach a numeric answer. Models are prompted with four chain-of-thought exemplars, and the final prediction is deemed correct only if the predicted numeric answer exactly matches the gold label. We use accuracy as the evaluation metric.
PIQA Bisk et al. (2020) is a benchmark for physical commonsense reasoning. Each of its 1,838 examples presents a naturalistic physical situation and two possible solutions, requiring models to choose the more plausible one. It evaluates a model’s ability to reason about everyday physical interactions. We use exact match as the evaluation metric.
SAMSum Gliwa et al. (2019) is a dialogue summarization dataset composed of multi-turn conversations between fictitious participants. The task requires models to generate concise and coherent summaries that capture key information. No in-context examples are used. We use rouge-1 score as the evaluation metric.
TriviaQA Joshi et al. (2017) contains 11,313 open-domain factoid questions authored by trivia enthusiasts, paired with Wikipedia-based ground-truth answers. Following prior work, each example includes five in-context QA exemplars, and model accuracy is measured by exact match between predicted and reference answers. We use exact match as the evaluation metric.
NQ Kwiatkowski et al. (2019) consists of 3,610 real anonymized Google search queries paired with short answers from Wikipedia articles. We follow prior work using 5-shot in-context prompting. We use exact match as the evaluation metric.
MMLU Hendrycks et al. (2020) (Massive Multitask Language Understanding) is a 57-subject multiple-choice benchmark spanning STEM, humanities, and social sciences. Each question has four answer choices, and models are evaluated under 5-shot settings with 5,000 test examples. We use exact match as the evaluation metric.
Appendix D Limitations
CoRE requires access to token‐level logits of ensembled LLMs to enforce consistency at both token and model level, preventing its use with closed‐source or black‐box LLM APIs. Moreover, determining when to ensemble remains an open question: if the main model is already confident or the assistant model exhibits low confidence, skipping ensembling may prevent unreliable outputs from degrading performance. Finally, identifying which models to ensemble is also worth exploring, while our model weights provide a soft balancing mechanism, future work could study more principled criteria for selecting beneficial model combinations.
Appendix E Potential Risks
The proposed CoRE, while enhancing robustness and reliability through token- and model-level consistency evaluation, also introduce potential risks that warrant careful consideration. CoRE relies on access to fine-grained token probability distributions from individual models to compute consistency scores, which limits its applicability to open-weight or transparent systems. When applied to closed-source or API-based models, this requirement may lead to implementation incompatibility or potential breaches of usage policies.
Appendix F Use Or Create Scientific Artifacts
Our work is built on public benchmarks and contributes new code resources to the community. For evaluation, we use widely adopted datasets including GSM8K, PIQA, SAMSum, TriviaQA, NQ, and MMLU, without making any changes to their original content. In addition, we introduce and release the codebase of CoRE (upon publication), providing a clear and well-structured repository to improve the accessibility of our research.
F.1 Cite Creators Of Artifacts
All external artifacts are properly credited to their original publications and repositories.
All benchmarks used in this work, including GSM8K Cobbe et al. (2021), PIQA Bisk et al. (2020), SAMSum Gliwa et al. (2019), TriviaQA Joshi et al. (2017), NaturalQuestions Kwiatkowski et al. (2019), and MMLU Hendrycks et al. (2009), are credited to their respective authors.
Each LLM model used in this work, including Llama-3-8B-Instruct Dubey et al. (2024), Mistral-7B-Instruct-v0.1 Jiang et al. (2023a), Qwen2.5-3b-Instruct Team (2024), InternLM2.5-7b-Chat Team (2023) and openchat-3.5-0106 Wang et al. (2023), is referenced through its official technical report or HuggingFace model card to ensure appropriate acknowledgment of all upstream contributions.
F.2 Discuss The License For Artifacts
We comply with the licenses of all artifacts used or released in this work. The benchmarks are distributed under Creative Commons or public-domain terms, following the conditions specified by their original maintainers.
Model checkpoints retain their original open-source licenses, and our own code and generated data are released under the MIT license. All license files are included in our repository and described in the supplementary README.
F.3 Artifact Use Consistent With Intended Use
We confirm that our use of datasets and pre-trained models is consistent with their intended purpose. Benchmarks are used solely for inference and evaluation, which aligns with their terms of service, and no modified model weights are redistributed. The released code is for inference only and does not support fine-tuning or commercial redistribution of the original checkpoints.
F.4 Data Contains Personally Identifying Info Or Offensive Content
Although the benchmarks may contain public text with personal names or informal language, we ensure responsible use by:
  • Using only passages that are part of the officially released benchmark datasets.
  • Applying keyword-based filtering to remove or mask offensive content in analysis outputs.
F.5 Documentation Of Artifacts
We provide detailed documentation of all evaluation datasets and model combinations used in our experiments. Specifically, Appendix C of the appendix describes the coverage, task type, and evaluation metric for each dataset (e.g., reasoning, summarization, and knowledge QA), while Appendix H the LLMs included in the ensemble along with their sizes and sources. Together, these materials ensure transparency regarding the domains, data characteristics, and model diversity involved in our study.
Appendix G Statistics For Data
Dataset statistics are summarize as follow:
G.1 GSM8K (Arithmetic Reasoning)
  • Number of entries: 1319
  • Average question length: 239.9 characters
  • Average answer length: 272.3 characters
G.2 PIQA (Commonsense Reasoning)
  • Number of entries: 1838
  • Average question length: 36.1 characters
  • Average answer length: 1.0 character
G.3 SAMSum (Summarization)
  • Number of entries: 819
  • Average question length: 521.6 characters
  • Average answer length: 108.8 characters
G.4 TriviaQA (Knowledge)
  • Number of entries: 6000
  • Average question length: 78.8 characters
  • Average answer length: 267.9 characters
G.5 NaturalQuestions (Knowledge)
  • Number of entries: 3610
  • Average question length: 47.7 characters
  • Average answer length: 24.7 characters
G.6 MMLU (Comprehensive Examination)
  • Number of entries: 14042
  • Average question length: 274.5 characters
  • Average answer length: 1.0 character
Appendix H Computational Experiments
All computational experiments in this work are fully reproducible, with details provided in the main text and the Appendix.
H.1 Model Size And Budget
For each LLM used, we specify its total parameter count as follows: Llama-3-8B-Instruct contains around 8 billion parameters; Mistral-7B-Instruct-v0.1, InternLM2.5-7b-Chat, and openchat-3.5-0106 contains around 7 billion parameters; Qwen2.5-3b-Instruct contains around 3 billion parameters. The total compute budget for all experiments is approximately 500 A100 GPU hours.
H.2 Experimental Setup And Hyperparameters
We describe experimental settings for all experiments in Section 4.1. For hyperparameter settings, we set $\sigma=0.5$ in the RBF kernel, $\alpha=1.0,\beta=1.0$ in the power function, and $\gamma=1.0$ in the sigmoid function. We clip the normalized model weight for the main LLM to be at least 0.5 to ensure its major contribution to the ensemble results.
H.3 Descriptive Statistics
For each result in the main text and Appendix, we report the mean across multiple runs.
H.4 Parameters For Packages
The existing packages used are specified as follows. We use PyTorch (v2.8.0) as the core deep learning framework, together with HuggingFace Transformers (v4.56.1) and Tokenizers (v0.22.0) for model implementation and text pre-processing. For efficient multi‑GPU and distributed training we rely on Accelerate (v1.10.1).
Appendix I Ai Assistants In Research Or Writing
In this paper, AI assistant tool is used to edit and improve the quality of the text, including checking the spelling, grammar, punctuation and clarity.